4). Angles For the given triangle, provide numerical values for sin(0), cos(®), and tan(o). Show (at least) major steps 45
Answer #1

Solution)

On applying Pythagoras Theorem

15^2=6^2 + A^2

So, A=13.74 m

Now,

Sin theta= perpendicular/hypotnuse=6/15 (Ans)

Cos theta=base/ hypotnuse=13.74/15 (Ans)

Tan theta=perpendicular/base=6/13.74 (Ans)

=====

Good luck!:)